(3 pts) Sue has $30,000 to use as a down payment on a house and can afford to pay $1000 per month for a mortgage. If the interest rate on a 15-year mortgage is 4.2% (this is an APR) compounded monthly. What is the highest price house she can afford using a 15-year mortgage? 2) s. (2 pts) Larry would like to retire in 20 years. He currently has $300,000 in his retirement account and is planning on depositing an additional $800 each month in his retirement account If Larry can earn an average APR of 7.2% per year compounded monthly, how much money will be in his retirement account in 20 years? 3) s. 4) (2 pts) You want to set up an endowed scholarship that will pay $6000 a year in perpetuity. If the appropriate interest rate is 4.5 % compounded annually, how much will you have to donate today to set up the endowed scholarship? 5) (2 pts) You want to buy a $25,000 car and have saved up $3,000 to use as a down payment. If you finance the rest at 4.5% (APR) compounded monthly for 5 years, what will your monthly payment be?
Answer #1

(2) Maximum Possible Monthly Repayments = $ 1000, Tenure of Mortgage = 15 years or 180 months and APR = 4.2 % per annum or (4.2/12) = 0.35 % per month

Maximum Possible Mortgage Amount = 1000 x (1/0.0035) x [1-{1/(1.0035)^(180)}] = $ 133377.73

Maximum Possible Down Payment = $ 30000

Maximum Possible House Price = Maximum Down payment + Maximum Mortgage = 30000 + 133377.73 = $ 163377.73

NOTE: Please raise separate queries for solutions to the remaining unrelated questions, as one query is limited to the solution of only one complete question (with a maximum of four sub-parts)